OpenStudy (anonymous):

help plz y=x^2-2x-2 y=4x+5 A (-1,1) and (-7,-23) B (-1,1) and (7,33) C (-1,33) and (7,1)

OpenStudy (anonymous):

D. No solution

OpenStudy (welshfella):

as they are both y= we can equate the expressions in x so x^2 - 2x - 2 = 4x + 5 x^2 - 6x - 7 = 0 can you solve this ?

OpenStudy (welshfella):

this will factor to the form (x + a)(x + b) = 0 you need to find a and b

OpenStudy (welshfella):

to find a and b you need to find 2 numbers whose product = -7 and whose sum is -6

OpenStudy (welshfella):

ab = -7 a + b = -6

OpenStudy (anonymous):

I know there two solution

OpenStudy (anonymous):

a=1, b=-6 c=-7

OpenStudy (welshfella):

the a and b are not the coefficients ax^2 + bx + c . Sorry if i have confused you . I could have used e ,f j and k any letters.

OpenStudy (welshfella):

the product is -7 * 1 = -7 and sum is - 7+ 1 = -6 so the equatiion becomes (x - 7)(x + 1) = 0 so what are the 2 values of x?

OpenStudy (anonymous):

x=7 1 x=-1 2

OpenStudy (welshfella):

yes x = 7 and x = -1 now you have to find the values of y for each value of x.

OpenStudy (welshfella):

do this by plugging in these values into y = 4x + 5 so for x = -1 y = 4(-1)

OpenStudy (welshfella):

y = 4(-1) + 5

OpenStudy (anonymous):

y=4(8)+5=33 y=4(-1)+5=1

OpenStudy (welshfella):

yes they are correct

OpenStudy (anonymous):

so the answer is (-1,1) and (7,33)?

OpenStudy (welshfella):

yes B

OpenStudy (anonymous):

thank you

OpenStudy (welshfella):

yw